# Beyond Redundancy: Toward Agile Resilience in Optical Networks to Overcome Unpredictable Disasters

### v1

@@ IV Validation through field evaluations
We successfully provisioned optical paths between two sites approximately 300 km apart and performed a database migration over the established optical link. The procedure was structured as follows:
@@
    Initial measurements were conducted to assess the performance characteristics of the transceivers at both endpoints.
@@
    • Measurement Phase (20 minutes): The received signals are captured at the receiver transponder.
    • Analysis Phase (40 minutes): The collected data was analyzed to extract key fiber characteristics such as attenuation, dispersion, and non-linearities.
@@
    • Measurement Phase (2.5 hours): The OLS was measured on different operating conditions to estimate the transmission physical parameters of the link.
    • Analysis Phase (1 hour): The system parameters were estimated and the amplifier configurations were optimized based on the calibration results to ensure stable and high-quality transmission.
@@
    Transponder settings were designed to define the corresponding lightpath according to the line system QoT estimation.
@@
Figure 3: Calibration results for the OLS from Dublin to Galway. (a) Computed optimal launch power before booster. (b) Computed accumulated GSNR after each EDFA. The optimal launch power is computed to have the highest flatten GSNR spectrum at the end of the line.
@@
Figure 4: Analyzed power profile with DLM. The vertical axis and horizontal axis show absolute power and distance from the Galway site, respectively. The black solid line and the red dotted line show before/after optimization.
Figure 5: Received WDM spectrum during validation, consisting of DLM 800G (130 GBd), 400G (63.1 GBd) and ASE noise loading (50 GHz) channels.
@@
After performing the parameter calibration of the physical model [12],the model was used to design the optimal amplifier configurations, including gain and tilt settings, and the optimal WDM channel launch power before the booster. The optimization target is the generalized signal-to-noise ratio (GSNR), which allows us to relate the optical transmission performance of the system, or QoT, to transceiver BER [13]. The design simulation results of the launch power optimization are reported in Fig. 3, illustrating the per-channel launch power configuration and the GSNR spectrum evolution after each span of the long-haul link. The GNSR was flatten along the full C-band spectrum to ensure QoT uniformity for all lightpaths crossing the link, assuming a scenario of meshed network or data center exchange [10]. Fig. 4 shows the power profile along the link from Galway to Dublin obtained with DLM, comparing the results before and after OLS calibration (black and red dotted lines, respectively). This tool enables visualization of the channel power evolution, allowing the detection of irregular losses within fiber spans and the localization of power profile variations along the spatial coordinate over time.
@@
The QoT of the long-haul link was validated after provisioning two 800G and four 400G channels and measuring the corresponding pre-FEC BER. Figure 5 reports the measured WDM spectrum after lightpath deployment. The difference between the estimated QoT at the design stage and the measured QoT after path establishment was conservative with a margin of roughly 1 dB. The results showed that the entire process, including system characterization, path design and establishment and database migration, was completed in approximately 6 hours. The migration was completed without errors, confirming the stability and performance of the optical link.
@@ VI Conclusion
Given that recent environmental shifts have amplified disaster unpredictability, the effectiveness of traditional redundancy-based approaches is failing. To address this, we proposed agile resilience, which enables dynamic resource reallocation across domains and operators. We used a field-deployed system to demonstrate characterization, path provisioning, and service migration within 6 hours, confirming our proposal’s feasibility. Future work includes AI-driven automation and satellite-based optics, making agile resilience a critical foundation for a sustainable information society.
@@ Acknowledgments
Toru Mano is a research engineer at NTT Network Innovation Laboratories. He received his B.E. and M.E. degrees from the University of Tokyo in 2009 and 2011, respectively. He joined NTT Laboratories in 2011. He received his Ph.D. degree in computer science and information technology from Hokkaido University in 2020. His research interests include network architectures, network optimization, and network softwarization.
Hideki Nishizawa is a senior research engineer at NTT Network Innovation Laboratories. He received B.E. and M.E. degrees in physics from Chiba University, Chiba, Japan, in 1994 and 1996, respectively. In 1996, he joined NTT Laboratories, Japan, where he has been engaged in research on open and disaggregated optical systems. He received the Ph.D. degree in information science and technology from Hokkaido University in 2025.
Takeo Sasai received the B.E., M.E. and Ph.D. degrees from the University of Tokyo, Japan, in 2015, 2017, and 2025 respectively. In 2017, he joined NTT Laboratories, Yokosuka, Japan. His research interests include physical layer monitoring, digital signal processing, and nonlinear modeling for optical fiber communication.
@@
Dmitrii Briantcev is a research fellow at the CONNECT Centre, Trinity College Dublin. He received the B.Sc. in Radiophysics from Saint Petersburg State University (2018) and the M.Sc. (2020) and Ph.D. (2023) in Electrical and Computer Engineering from KAUST, Saudi Arabia. He works on optical communications and digital twins, integrating simulation, data-driven channel modeling, and testbed validation.
@@
Devika Dass is a research fellow at the CONNECT Centre, Trinity College Dublin. Her research interest encompasses My research interests include exploring higher-order data modulation for a converged optical metro-access network, analog Radio-over-Fiber and Free-space optics. She earned her Ph.D. in 2023 from Dublin City University, master’s in Communication Engineering from VIT University, Vellore, India (2017) and undergraduate in Electronics and Telecommunication from Amity University, India (2014).
@@
Brandt Bashaw is a Ph.D student at the CONNECT Centre, Trinity College Dublin. He received his B.Sc. in Electrical Engineering from Brigham Young University in 2023. His research interests include classical optical communications, quantum optics, and quantum computing.
@@
Eoin Kenny is the Innovation, Research and Development manager at HEAnet (Ireland’s National Education and Research Network). His research interests are in the areas of advanced network architectures, quantum communications, distribution of timing services, fibre sensing, and control and management planes for future networks. He received a degree in Electronic Engineering from University College Dublin in 1993.
@@
Marco Ruffini is full professor at Trinity College Dublin, leading the OpenIreland lab and the Optical and Wireless networks research group. His main research is in the area of Intelligent networks, where he carries out pioneering work on the convergence of fixed-mobile and access-metro networks, AI-based control and digital twins of optical networks, quantum networking and fibre sensing. He has been invited to share his vision through several keynote and talks at major international conferences across the world. He authored over 220 international publications, 10 patents, contributed to industry standards and secured research funding for over € 14 milion, and contributed his novel virtual Dynamic Bandwidth Allocation (vDBA) concept to the BroadBand Forum standardisation body.
@@
Yoshiaki Sone is a senior research engineer at NTT Network Innovation Laboratories. He received his M.E. degree in electronics engineering from Tohoku University, Miyagi, in 2003. In 2003, he joined NTT Laboratories to focus his research on network engineering technologies for optical transport networks.
@@
Koichi Takasugi received the B.E. degree in computer science from Tokyo Institute of Technology in1995, the M.E. degree from Japan Advanced Institute of Science and Technology in 1997 and Ph.D. in engineering from Waseda University in 2004. In 1997, he joined NTT Laboratories, Japan, where he has been engaged in research on wireless and optical transport networks.
@@
Daniel Kilper is Professor of Future Communication Networks and SFI CONNECT Centre Director at Trinity College Dublin, Ireland. He holds an adjunct faculty appointment at Columbia University Data Science Institute and College of Optical Sciences, University of Arizona. He is CTO and co-founder of Palo Verde Networks, Inc., and is the Green Internet and Service Provisioning topical area editor for IEEE Transactions on Green Communications and Networking. He co-chairs the IEEE International Network Generations Roadmap (INGR) Optics Working Group. He received M.S. (1992) and Ph.D. (1996) degrees in physics from the University of Michigan. From 2000 to 2013, he was a member of the technical staff at Bell Labs. His research is aimed at solving fundamental and real-world problems in communication networks, addressing interdisciplinary challenges for smart cities, sustainability, and digital equity.

### v2

@@ IV-B Procedure and time record
We successfully provisioned optical paths between two sites approximately 300 km apart and performed a database migration over the established optical link. The experimental procedure was as follows:
@@
    The transceivers’ performance characteristics were assessed at both endpoints.
@@
    • Measurement (20 minutes): The received signals are captured at the receiver transponder.
    • Analysis (40 minutes): The collected data was analyzed to extract key fiber characteristics such as attenuation, dispersion, and nonlinearities.
@@
    • Measurement (2.5 hours): The OLS was measured under different operating conditions to estimate the transmission physical parameters of the link.
    • Analysis (1 hour): The system parameters were estimated and the amplifier configurations were optimized based on the calibration results to ensure stable and high-quality transmission.
@@
    Transponder settings were designed to define the corresponding optical path according to the line system QoT estimation.
@@
Figure 3: Calibration results for the OLS from Dublin to Galway. (a) Computed optimal launch power before booster. (b) Computed accumulated GSNR after each EDFA. The optimal launch power is computed to have the highest flattened GSNR spectrum at the end of the line.
(b)Refer to caption
@@
Figure 4: DLM validation. (a) Analyzed power profile. The vertical axis and horizontal axis show absolute power and distance from the Galway site, respectively. The blue solid line and the red dotted line show before/after optimization. (b) Received WDM spectrum during validation, consisting of DLM 800G (130 GBd), 400G (63.1 GBd) and ASE noise loading (50 GHz) channels.
@@
After performing the parameter calibration of the physical model [11], the model was used to design the optimal amplifier configurations, including gain and tilt settings, and the optimal WDM channel launch power before the booster. The optimization target is the generalized signal-to-noise ratio (GSNR), which allows us to relate the optical transmission performance of the system, or QoT, to transceiver BER [12]. The design simulation results of the launch power optimization are reported in Fig. 3, illustrating the per-channel launch power configuration and the GSNR spectrum evolution after each span of the long-haul link. The GSNR was flattened along the full C-band spectrum to ensure QoT uniformity for all optical paths crossing the link, assuming a scenario of meshed network or data center exchange [9]. Fig. 4 (a) shows the power profile along the link from Galway to Dublin obtained with DLM, comparing the results before and after OLS calibration (black and red dotted lines, respectively). This tool enables visualization of the channel power evolution, allowing the detection of irregular losses within fiber spans and the localization of power profile variations along the spatial coordinate over time.
IV-C Results and Discussions
@@
The QoT of the long-haul link was validated after provisioning two 800G and four 400G channels and measuring the corresponding pre-FEC BER. Figure 4 (b) reports the measured WDM spectrum after optical path deployment. The difference between the estimated QoT at the design stage and the measured QoT after path establishment was conservative with a margin of roughly 1 dB. The results showed that the entire process, including system characterization, path design and establishment and database migration, was completed in approximately 6 hours. The migration was completed without errors, confirming the stability and performance of the optical link.
While the 6-hour provisioning time is faster than manual configuration, there is an inherent trade-off between provisioning speed and optimization depth. The optimal balance in disaster recovery depends on factors such as link conditions, operational urgency, and specific requirements from users or authorities. Simpler optimization methods could save several hours but may result in suboptimal configurations. In particular, they may not adequately cope with nonlinear optical effects or unexpected loss variations in newly constructed lines using unknown dark fibers, thereby posing risks of link establishment failure or severe instability. This paper prioritizes the reliability of service restoration, adopting an approach that combines full characterization and in-depth optimization realized by DLM and OLS calibration. Nevertheless, further reducing provisioning time for scenarios where speed is the primary concern remains an important topic for future work.
Note that, as the temporary line system is newly constructed specifically for disaster recovery and carries no pre-existing traffic, the OLS calibration can be executed without causing any service disruption.
@@ VI Conclusion
Given that recent environmental shifts have amplified disaster unpredictability, the effectiveness of traditional redundancy-based approaches is failing. To address this, we proposed agile resilience, which enables dynamic resource reallocation across domains and operators. We used a field-deployed system to demonstrate characterization, path provisioning, and service migration within 6 hours, confirming our proposal’s feasibility. Future work includes AI-driven automation and satellite-based optics, positioning agile resilience as a critical foundation for a sustainable information society.
@@ Acknowledgments
Toru Mano is a research engineer at NTT Network Innovation Laboratories. He received his B.E. and M.E. degrees from the University of Tokyo in 2009 and 2011, respectively. He joined NTT Laboratories in 2011. He received his Ph.D. degree in computer science and information technology from Hokkaido University in 2020.
Hideki Nishizawa is a Senior Research Engineer at NTT Network Innovation Laboratories, Japan. He received B.E. and M.E. degrees in Physics from Chiba University in 1994 and 1996, respectively, and a Ph.D. in Information Science and Technology from Hokkaido University in 2025.
Takeo Sasai received the B.E., M.E. and Ph.D. degrees from the University of Tokyo, Japan, in 2015, 2017, and 2025 respectively. In 2017, he joined NTT Laboratories, Yokosuka, Japan.
@@
Dmitrii Briantcev is a research fellow at the CONNECT Centre, Trinity College Dublin. He received the B.Sc. in Radiophysics from Saint Petersburg State University (2018) and the M.Sc. (2020) and Ph.D. (2023) in Electrical and Computer Engineering from KAUST, Saudi Arabia.
@@
Devika Dass is a research fellow at the CONNECT Centre, Trinity College Dublin. She earned her Ph.D. in 2023 from Dublin City University, master’s in Communication Engineering from VIT University, Vellore, India (2017) and undergraduate in Electronics and Telecommunication from Amity University, India (2014).
@@
Brandt Bashaw is a Ph.D. student at the CONNECT Centre, Trinity College Dublin. He received his B.Sc. in Electrical Engineering from Brigham Young University in 2023. His research interests include classical optical communications, quantum optics, and quantum computing.
@@
Eoin Kenny is the Innovation, Research and Development manager at HEAnet (Ireland’s National Education and Research Network). He received a degree in Electronic Engineering from University College Dublin in 1993.
@@
Marco Ruffini is full professor at Trinity College Dublin, leading the OpenIreland lab and the Optical and Wireless networks research group. He authored over 220 international publications, and secured research funding for over € 14 milion, and contributed his novel virtual Dynamic Bandwidth Allocation (vDBA) concept to the BroadBand Forum standardisation body.
@@
Yoshiaki Sone is a senior research engineer at NTT Network Innovation Laboratories. He received his M.E. degree in electronics engineering from Tohoku University, Miyagi, in 2003. In 2003, he joined NTT Laboratories.
@@
Koichi Takasugi received the B.E. degree in computer science from Tokyo Institute of Technology in 1995, the M.E. degree from Japan Advanced Institute of Science and Technology in 1997 and Ph.D. in engineering from Waseda University in 2004. In 1997, he joined NTT Laboratories, Japan.
@@
Daniel Kilper is Professor of Future Communication Networks and SFI CONNECT Centre Director at Trinity College Dublin, Ireland. He received M.S. (1992) and Ph.D. (1996) degrees in physics from the University of Michigan. His research is aimed at solving fundamental and real-world problems in communication networks, addressing interdisciplinary challenges for smart cities, sustainability, and digital equity.